# Martian time-series unraveled: A multi-scale nested approach with factorial variational autoencoders

### v4

@@ VI Discussion and Conclusion
Future work involves letting the mother wavelet in the scattering network to be learned, e.g., using [69]. In that setting, the mother wavelet is no longer specified a priori, but is rather parametrized by a few learnable parameters. Those parameters can then be learned to adapt the shape of the mother wavelet–and thus of all the signal statistics that the scattering network computes–to better minimize the loss at hand. Because the parametrization only adds a few parameters, overfitting is not an issue, even in noisy settings, as demonstrated in [53].
Appendix A All identified clusters and their representative waveforms
@@
In a similar study, [64] also aimed at detecting and clustering microevents and noise patterns in the seismic data recorded during the InSight mission on Mars. The authors identified nine main clusters of events and noise patterns in the seismic data when focusing on $100\,\mathrm{s}$ time windows. Similar to our observations with the finest timescale used in our results, [64] mostly identified transient events, such as pressure drops and various types of glitches. Since their timescale was slightly longer than our finest timescale, they also discovered repeating sequences of glitches with relatively constant time offsets. This is analogous to the two clusters we identify in the $3.4$-minute timescale presented in Figs. 11(a) and 11(f), whose time histograms also align with the typical occurrence time of glitches (see Figs. 10(a) and 10(f)). Regarding pressure drops, [64] reported that these were mainly spread across three clusters. Although the aligned waveforms of clusters 0, 7, and 8 identified by our approach (see Figs. 9(a), 9(h) and 9(i)) do not indicate pressure drops, the latent space visualization of pressure drops from the catalog at the finest scale (see Fig. 7(b)) shows that these pressure drops are mostly concentrated in cluster 7, with the rest primarily in clusters 0 and 8. Additionally, [64] characterized daily and seasonal variations in event occurrence related to temperature changes, which we similarly observed in Fig. 4.
@@
Appendix B fVAE decoder reconstruction quality
@@
Appendix C Unsupervised separation of Marsquakes
@@
Next, we use these time windows as representative samples from the background noise ($\mathbf{s}_{i}$ in the context of 10) with the goal of separating the clean marsquake signal. We solve the optimization problem in 10 with 1000 L-BFGS iterations. The results are depicted in Fig. 20. There are clearly three glitches that we have successfully separated, along with the background noise. Through this example, we demonstrated the effectiveness of our multi-scale source separation method—based on prior information obtained from the fVAE—when the source (the marsquake in this case) is not a prominent source in the entire dataset, i.e., it is an outlier.
Appendix D Stylized example
We present a multi-scale synthetic dataset (see Fig. 21) used to validate our model using ground truth sources. We first describe how the dataset is constructed and then we present the results of our clustering and source separation algorithm when applied to this dataset.
D-A Synthetic multi-scale dataset
@@
  • Large scale ($w=2^{12}$). In this timescale, the signal alternates between two different noises to mimic the change in background noise due to day-night periodicity. The first noise is a white noise $\mathbf{x}_{1}(t)$ and the second one is a multifractal random noise $\mathbf{x}_{2}(t)$ which is a non-Gaussian noise with intermittency [72]. The large-scale signal is then $\mathbf{x}_{\text{large}}(t)=(1-\theta(t))\,\mathbf{x}_{1}(t)+\theta(t)\,\mathbf{x}_{2}(t)$ where $\theta(t)$ is equal to 00 on $[0,w]$ and $1$ on $[w+\eta,2w-\eta]$ for small $\eta$, is $2w$-periodic and has a smooth junction at $t=w$ and at $t=2w$.
@@
The resulting dataset is a long time-series $\mathbf{x}=\mathbf{x}_{\text{large}}+\mathbf{x}_{\text{medium}}+\mathbf{x}_{\text{fine}}$.
D-B Synthetic data clustering
While we are aware of the underlying sources that create this data, we choose to make an uninformed decision in regards to the number of clusters per scale to mimic a more realistic setting. To this end, we use the same architecture and same hyperparameters as the fVAE trained in the context of clustering and source separation data form Mars, with the distinction that we chose our timescales for scattering covariance generation based on the duration of sources in our multi-scale dataset.

### v5

@@ VI Discussion and Conclusion
Future work involves letting the mother wavelet in the scattering network to be learned, e.g., using [71]. In that setting, the mother wavelet is no longer specified a priori, but is rather parametrized by a few learnable parameters. Those parameters can then be learned to adapt the shape of the mother wavelet–and thus of all the signal statistics that the scattering network computes–to better minimize the loss at hand. Because the parametrization only adds a few parameters, overfitting is not an issue, even in noisy settings, as demonstrated in [53]. Additionally, training the fVAE in the embedding space of pretrained seismic foundation models [72] could leverage rich representations from large-scale seismic datasets to further enhance performance.
Appendix A Algorithm overview: Step-by-step process
Below we provide a detailed walkthrough of the proposed unsupervised multi-scale clustering and source separation framework, following the flow of Fig. 1 from left to right. This description enumerates the main steps required to apply our approach.
  1. 1.
    Pyramidal wavelet scattering spectra computation
    • Begin with the raw multichannel time series dataset, denoted as $\mathbf{x}(t)$.
    • Transform the data into the pyramidal wavelet scattering spectra representation, $\Psi(\mathbf{x})$, as described in Section II and 3. This step captures the multi-scale, non-Gaussian characteristics of the data by computing scattering features over multiple window sizes (pyramidal averaging).
    • The result is a set of multi-scale feature vectors that summarize the signal’s structure across a wide range of timescales (see multi-scale feature extractor in Fig. 1).
  2. 2.
    Probabilistic multi-scale clustering
    • Feed the multi-scale representations into the fVAE, which is trained to model the joint distribution of features at all timescales (see Section III and 9).
    • The fVAE simultaneously performs probabilistic clustering at each timescale, assigning each data window to a cluster that represents a prominent source or process in the dataset.
    • The model learns both the cluster assignments and the underlying distribution of features for each cluster, which will serve as prior information for separation (see fVAE module in Fig. 1).
  3. 3.
    Identification of prominent source clusters
    • Analyze the clustering results to identify clusters corresponding to sources of interest (e.g., glitches, wind bursts, background noise).
    • This step may involve inspecting cluster occurrence histograms and representative waveforms, as shown in Section V-D.
  4. 4.
    Formulation of the source separation problem
    • For a target source (cluster), collect representative samples from the identified cluster(s) to serve as prior information.
    • Formulate the source separation as an optimization problem in the scattering spectra space, as described in Section IV. The goal is to recover the target source. The objective combines three loss terms:
      • Prior consistency ($\mathcal{L}_{\text{prior}}$), ensuring the separated source matches the cluster statistics (cf. 10a),
      • Cross-independence ($\mathcal{L}_{\text{cross}}$), promoting independence between the separated source and residual (cf. 10b),
      • Data fidelity ($\mathcal{L}_{\text{data}}$), ensuring the sum of separated sources reconstructs the observed mixture (cf. 10c).
    • See the source separation module in Fig. 1 for a high-level depiction of this step.
  5. 5.
    Optimization and recovery of source signals
    • Solve the combined optimization problem, as presented in $1$, to recover the target source in the time domain, using the learned priors and constraints from the previous step.
    • The optimization is performed for each window or segment, and can be parallelized (final stage in Fig. 1).
Appendix B Derivation of the fVAE loss function
To overcome the intractability of minimizing the negative log-likelihood, i.e., minimizing the expected value of $\log p_{\theta}(\mathbf{u})$, we instead minimize the negative lower bound (ELBO). Here, we detail the derivation of the objective function presented in 9. The derivation proceeds as follows:
  • The derivation begins in 12a with the original objective of minimizing the expected negative log-likelihood of the data under the model. In 12b, latent variables $\mathbf{z},\mathbf{y}$ are introduced via a variational distribution $q_{\phi}(\mathbf{z},\mathbf{y}|\mathbf{u})$, enabling a tractable approximation through sampling.
  • Line 12c rewrites the marginal likelihood using Bayes’ rule, expressing $\log p_{\theta}(\mathbf{u})$ as a ratio of the joint and conditional distributions. In 12d, an identity involving the variational distribution is inserted by multiplying and dividing by $q_{\phi}(\mathbf{z},\mathbf{y}|\mathbf{u})$, allowing for further decomposition. This is rearranged in 12e to separate model and variational terms. 12f then expands the log expression, revealing the reverse KL divergence between $q_{\phi}(\mathbf{z},\mathbf{y}|\mathbf{u})$ and $p_{\theta}(\mathbf{z},\mathbf{y}|\mathbf{u})$. The inequality in 12g follows by dropping this non-negative KL term, yielding a variational lower bound (ELBO) on the negative log-likelihood.
  • In 12h, we substitute the generative model factorization from 4 and the inference model factorization from 6 into the joint and variational distributions.
  • Line 12i converts the logarithm of the product into a sum of logarithms. In 12k, we interchange the expectation and summation operators, and factor the expectation over the categorical and Gaussian latent variables.
  • Finally, 12l through 12n decompose this lower bound into three interpretable components by expanding the logarithmic expression and regrouping terms: a per-scale reconstruction loss (ensuring the decoder can reconstruct the input scattering spectra), a categorical KL divergence for $y_{i}$ (regularizing the cluster assignment probabilities), and a Gaussian mixture KL divergence for $z_{i}$ (regularizing the latent representations within each cluster), resulting in the final form of the objective function presented in the main text.
$=$ $\min_{\boldsymbol{\theta},\boldsymbol{\phi}}\operatorname*{\mathbb{E}}_{\begin{subarray}{c}\mathbf{u}\sim p(\mathbf{u})\\[-1.0pt] \mathbf{z},\mathbf{y}\sim q_{\phi}(\mathbf{z},\mathbf{y}|\mathbf{u})\end{subarray}}\big[-\log p_{\theta}(\mathbf{u})\big]$ (12b)
$=$ $\min_{\boldsymbol{\theta},\boldsymbol{\phi}}\operatorname*{\mathbb{E}}_{\begin{subarray}{c}\mathbf{u}\sim p(\mathbf{u})\\[-1.0pt] \mathbf{z},\mathbf{y}\sim q_{\phi}(\mathbf{z},\mathbf{y}|\mathbf{u})\end{subarray}}\Big[-\log\frac{p_{\theta}(\mathbf{u},\mathbf{y},\mathbf{z})}{p_{\theta}(\mathbf{z},\mathbf{y}|\mathbf{u})}\Big]$ (12c)
$=$ $\min_{\boldsymbol{\theta},\boldsymbol{\phi}}\operatorname*{\mathbb{E}}_{\begin{subarray}{c}\mathbf{u}\sim p(\mathbf{u})\\[-1.0pt] \mathbf{z},\mathbf{y}\sim q_{\phi}(\mathbf{z},\mathbf{y}|\mathbf{u})\end{subarray}}\Big[-\log\frac{q_{\phi}(\mathbf{z},\mathbf{y}|\mathbf{u})}{q_{\phi}(\mathbf{z},\mathbf{y}|\mathbf{u})}\frac{p_{\theta}(\mathbf{u},\mathbf{y},\mathbf{z})}{p_{\theta}(\mathbf{z},\mathbf{y}|\mathbf{u})}\Big]$ (12d)
$=$ $\min_{\boldsymbol{\theta},\boldsymbol{\phi}}\operatorname*{\mathbb{E}}_{\begin{subarray}{c}\mathbf{u}\sim p(\mathbf{u})\\[-1.0pt] \mathbf{z},\mathbf{y}\sim q_{\phi}(\mathbf{z},\mathbf{y}|\mathbf{u})\end{subarray}}\Big[-\log\frac{q_{\phi}(\mathbf{z},\mathbf{y}|\mathbf{u})}{p_{\theta}(\mathbf{z},\mathbf{y}|\mathbf{u})}\frac{p_{\theta}(\mathbf{u},\mathbf{y},\mathbf{z})}{q_{\phi}(\mathbf{z},\mathbf{y}|\mathbf{u})}\Big]$ (12e)
$=$ $\min_{\boldsymbol{\theta},\boldsymbol{\phi}}\operatorname*{\mathbb{E}}_{\begin{subarray}{c}\mathbf{u}\sim p(\mathbf{u})\\[-1.0pt] \mathbf{z},\mathbf{y}\sim q_{\phi}(\mathbf{z},\mathbf{y}|\mathbf{u})\end{subarray}}\Big[\log\frac{p_{\theta}(\mathbf{z},\mathbf{y}|\mathbf{u})}{q_{\phi}(\mathbf{z},\mathbf{y}|\mathbf{u})}-\log\frac{p_{\theta}(\mathbf{u},\mathbf{y},\mathbf{z})}{q_{\phi}(\mathbf{z},\mathbf{y}|\mathbf{u})}\Big]$ (12f)
$\leq$ $\min_{\boldsymbol{\theta},\boldsymbol{\phi}}\operatorname*{\mathbb{E}}_{\begin{subarray}{c}\mathbf{u}\sim p(\mathbf{u})\\[-1.0pt] \mathbf{z},\mathbf{y}\sim q_{\phi}(\mathbf{z},\mathbf{y}|\mathbf{u})\end{subarray}}\Big[-\log\frac{p_{\theta}(\mathbf{u},\mathbf{y},\mathbf{z})}{q_{\phi}(\mathbf{z},\mathbf{y}|\mathbf{u})}\Big]$ (12g)
$=$ $\min_{\boldsymbol{\theta},\boldsymbol{\phi}}\operatorname*{\mathbb{E}}_{\begin{subarray}{c}\mathbf{u}\sim p(\mathbf{u})\\[-1.0pt] \mathbf{z},\mathbf{y}\sim q_{\phi}(\mathbf{z},\mathbf{y}|\mathbf{u})\end{subarray}}\Big[-\log\frac{\prod_{i=0}^{s-1}p_{\theta}(\mathbf{u}_{i}|\mathbf{z}_{i})\,p_{\theta}(\mathbf{z}_{i}|y_{i})\,p_{\theta}(y_{i})}{\prod_{i=0}^{s-1}q_{\phi}(\mathbf{z}_{i}|y_{i},\mathbf{u})\,q_{\phi}(y_{i}|\mathbf{u})}\Big]$ (12h)
$=$ $\min_{\boldsymbol{\theta},\boldsymbol{\phi}}\operatorname*{\mathbb{E}}_{\begin{subarray}{c}\mathbf{u}\sim p(\mathbf{u})\\[-1.0pt] \mathbf{z},\mathbf{y}\sim q_{\phi}(\mathbf{z},\mathbf{y}|\mathbf{u})\end{subarray}}\Big[-\sum_{i=0}^{s-1}\log\frac{p_{\theta}(\mathbf{u}_{i}|\mathbf{z}_{i})\,p_{\theta}(\mathbf{z}_{i}|y_{i})\,p_{\theta}(y_{i})}{q_{\phi}(\mathbf{z}_{i}|y_{i},\mathbf{u})\,q_{\phi}(y_{i}|\mathbf{u})}\Big]$ (12i)
$=$ $\min_{\boldsymbol{\theta},\boldsymbol{\phi}}\sum_{i=0}^{s-1}\operatorname*{\mathbb{E}}_{\mathbf{u}\sim p(\mathbf{u})}\,\Big[\operatorname*{\mathbb{E}}_{\begin{subarray}{c}y_{i}\sim q_{\phi}(y_{i}|\mathbf{u})\\[-1.0pt] \mathbf{z}_{i}\sim q_{\phi}(\mathbf{z}_{i}|y_{i},\mathbf{u})\end{subarray}}$ (12j)
$\quad\Big[-\log\frac{p_{\theta}(\mathbf{u}_{i}|\mathbf{z}_{i})\,p_{\theta}(\mathbf{z}_{i}|y_{i})\,p_{\theta}(y_{i})}{q_{\phi}(\mathbf{z}_{i}|y_{i},\mathbf{u})\,q_{\phi}(y_{i}|\mathbf{u})}\Big]\Big]$ (12k)
$=$ $\min_{\boldsymbol{\theta},\boldsymbol{\phi}}\sum_{i=0}^{s-1}\operatorname*{\mathbb{E}}_{\mathbf{u}\sim p(\mathbf{u})}\,\Big[\operatorname*{\mathbb{E}}_{\mathbf{z}_{i}\sim q_{\phi}(\mathbf{z}_{i}|y_{i},\mathbf{u})}\,\underbrace{\big[-\log p_{\theta}(\mathbf{u}_{i}|\mathbf{z}_{i})\big]}_{\begin{subarray}{c}\text{Per-scale reconstruction loss}\end{subarray}}$ (12l)
$\quad+\underbrace{\mathbb{KL}\,\big(q_{\phi}(y_{i}|\mathbf{u})\,||\,p_{\theta}(y_{i})\big)}_{\begin{subarray}{c}\text{Categorical prior on }y_{i}\end{subarray}}$ (12m)
$\quad+\operatorname*{\mathbb{E}}_{y_{i}\sim q_{\phi}(y_{i}|\mathbf{u})}\,\underbrace{\big[\mathbb{KL}\,\big(q_{\phi}(\mathbf{z}_{i}|y_{i},\mathbf{u})\,||\,p_{\theta}(\mathbf{z}_{i}|y_{i})\big)\big]}_{\begin{subarray}{c}\text{Gaussian mixture prior on }\mathbf{z}_{i}\end{subarray}}\Big]$ (12n)
Appendix C Optimal value of source separation objective function
As a quantitative measure of the source separation performance, we derive the optimal value of the objective function in 11. Assuming that the algorithm successfully separated the source $s_{1}$, one can derive a typical value for the loss $\mathcal{L}=\mathcal{L}_{\text{prior}}+\mathcal{L}_{\text{cross}}+\mathcal{L}_{\text{data}}$. This value can be used to quantify the separation in our algorithm.
Working term by term, the first term writes
$\mathcal{L}_{\text{prior}}(\mathbf{s}_{1})=\sum_{i=1}^{N}\frac{\Big\|\Psi_{k}\big(\mathbf{s}_{1}\big)-\Psi_{k}\big(\mathbf{s}^{i}_{1}\big)\Big\|_{2}^{2}}{\sigma^{2}\big(\Psi_{k}\big(\mathbf{s}^{i}_{1}\big)\big)}.$
In the best case, our retrieved statistics $\Psi(s_{1})$ have the same mean and the same variance as the observed population of statistics $\{\Psi(s^{i}_{1})\}_{i=1}^{N}$. Furthermore, we assume that the population of statistics $\{\Psi_{k}(s^{i}_{1})\}_{i=1}^{N}$ follows a multivariate independent Gaussian that is independent from $\Psi(s_{1})$. In that case, each summed term in $\mathcal{L}_{\text{prior}}(\mathbf{s}_{1})$ follows twice a $\chi^{2}$ distribution with $r$ degrees of freedom, where $r$ is the number of coefficients in $\Psi$. The term $\mathcal{L}_{\text{prior}}(\mathbf{s}_{1})$ is thus equal to $2Nr$ in expectation (or $2N$ if the $\|\cdot\|_{2}$ is normalized by $r$).
Regarding the third term
$\mathcal{L}_{\text{data}}(\mathbf{s}_{1})=\sum_{i=1}^{N}\frac{\Big\|\Psi_{k}\left(\mathbf{x}-\mathbf{s}_{1}+\mathbf{s}^{i}_{1}\right)-\Psi_{k}\left(\mathbf{x}\right)\Big\|^{2}_{2}}{\sigma^{2}\big(\Psi_{k}\big(\mathbf{x}+\mathbf{s}^{i}_{1}\big)\big)}$
if we assume that the variance of the coefficients $\sigma^{2}\big(\Psi_{k}\big(\mathbf{x}+\mathbf{s}^{i}_{1}\big)$ is close to the variance of the same coefficients evaluated on the mixture signal $\sigma^{2}\big(\Psi_{k}\big(\mathbf{x}\big)\big)$, then, one can derive similarly a typical value of $2Nr$ in case of successful separation.
The second term
$\mathcal{L}_{\text{cross}}(\mathbf{s}_{1})=\sum_{i=1}^{N}\frac{\Big\|\Psi_{k}\left(\mathbf{s}^{i}_{1},\mathbf{x}-\mathbf{s}_{1}\right)\Big\|^{2}_{2}}{\sigma^{2}\big(\Psi_{k}\big(\mathbf{s}^{i}_{1},\mathbf{x}\big)\big)},$
can be assumed to be negligible against the other ones in case of successful separation, where the terms $s_{1}^{i}$ and $x-s_{1}$ should be decorrelated.
In the case of successful separation, the loss $\mathcal{L}$ should thus be equal to $4Nr$ in expectation.
Appendix D All identified clusters and their representative waveforms
@@
In a similar study, [64] also aimed at detecting and clustering microevents and noise patterns in the seismic data recorded during the InSight mission on Mars. The authors identified nine main clusters of events and noise patterns in the seismic data when focusing on $100\,\mathrm{s}$ time windows. Similar to our observations with the finest timescale used in our results, [64] mostly identified transient events, such as pressure drops and various types of glitches. Since their timescale was slightly longer than our finest timescale, they also discovered repeating sequences of glitches with relatively constant time offsets. This is analogous to the two clusters we identify in the $3.4$-minute timescale presented in Figs. 11(a) and 11(f), whose time histograms also align with the typical occurrence time of glitches (see Figs. 10(a) and 10(f)).
Regarding pressure drops, [64] reported that these were mainly spread across three clusters. Although the aligned waveforms of clusters 0, 7, and 8 identified by our approach (see Figs. 9(a), 9(h) and 9(i)) do not indicate pressure drops, the latent space visualization of pressure drops from the catalog at the finest scale (see Fig. 7(b)) shows that these pressure drops are mostly concentrated in cluster 7, with the rest primarily in clusters 0 and 8. Additionally, [64] characterized daily and seasonal variations in event occurrence related to temperature changes, which we similarly observed in Fig. 4.
@@
Appendix E fVAE decoder reconstruction quality
@@
Appendix F Baseline glitch separation comparison
We have included a baseline comparison with the NASA InSight glitch separation algorithm [69], which represents the current state-of-practice for processing InSight seismic data. We illustrate the results of this comparison in Fig. 20.
This method employs a multi-step pipeline including data decimation, deconvolution with band-pass filtering, time derivative calculation to convert acceleration steps into impulse-like signals, and threshold-based detection followed by parametric modeling of glitch signatures. The separation process assumes observed data as a linear combination of glitch and background signal, then fits a three-parameter model (amplitude, offset, and linear trend) through nonlinear least squares optimization before subtracting the fitted glitch from the original data. While this approach represents established practice, it has fundamental limitations compared to our method. The NASA algorithm relies on explicit parametric modeling of glitch functional forms and assumes simple linear mixing, whereas our approach learns the statistical structure of background noise through wavelet scattering spectra without making assumptions about source morphology. As demonstrated in Fig. 20, our method achieves better separation quality by leveraging learned priors from clustering rather than rigid parametric templates, highlighting the advantage of data-driven approaches for complex planetary phenomena where source characteristics are not well understood a priori. Specifically, as confirmed by experts at the InSight team, our approach removes multiple glitches that the baseline has ignored while also performing better on glitches that have a spiking precursor, which the baseline method struggles to fully separate (compare third row of Fig. 5 and Fig. 20.)
Appendix G Comparison with ICA
To check whether the classical unsupervised source separation methods are able to perform source separation in our setting, we implemented ICA on our test cases including glitch separation and wind burst separation. As anticipated from our problem formulation, ICA failed to produce accurate separations for the desired sources. This failure confirms our initial assessment that traditional methods designed for stationary sources are inadequate for the complex temporal phenomena encountered in planetary seismic data. ICA’s poor performance stems from the fundamental mismatch between its core assumptions and the characteristics of Martian seismic data. The method operates on fixed time windows with stationarity assumptions, while our sources span vastly different timescales (51.2 seconds to 54.6 minutes) and exhibit strong non-stationarity through diurnal and seasonal variations as demonstrated in Fig. 4. Additionally, ICA relies solely on statistical independence without domain-specific knowledge, whereas our successful separations depend critically on learned priors from clustering similar source signatures. This comparison validates our multi-scale approach and demonstrates why domain-aware methodology is essential for analyzing complex planetary phenomena that systematically violate the assumptions underlying traditional signal processing methods.
We demonstrate ICA’s limitations through comprehensive comparison figures. Figure Fig. 21 shows ICA applied to glitch separation across three representative time windows. While ICA successfully decomposes the seismic data into independent components, none of the resulting components cleanly isolate the glitch signals. Instead, the glitches are distributed across multiple components mixed with background noise, making it impossible to achieve clean separation. Similarly, Figure Fig. 22 demonstrates ICA’s failure for wind burst separation, where the wind signatures are fragmented across components rather than being cleanly extracted. In contrast to our method’s targeted separation (shown in the main results), ICA’s blind decomposition produces components that lack physical interpretability and fail to align with the underlying seismic source mechanisms. These results quantitatively demonstrate the superiority of our domain-aware, multi-scale approach over traditional blind source separation methods.
Appendix H Unsupervised separation of Marsquakes
@@
Next, we use these time windows as representative samples from the background noise ($\mathbf{s}_{i}$ in the context of 11) with the goal of separating the clean marsquake signal. We solve the optimization problem in 11 with 1000 L-BFGS iterations and a relative termination tolerance of 0.1. The optimization took 40 minutes to complete. The results are depicted in Fig. 23. There are clearly three glitches that we have successfully separated, along with the background noise. Through this example, we demonstrated the effectiveness of our multi-scale source separation method—based on prior information obtained from the fVAE—when the source (the marsquake in this case) is not a prominent source in the entire dataset, i.e., it is an outlier.
Appendix I Stylized example
We present a multi-scale synthetic dataset used to validate our model using ground truth sources. We first describe how the dataset is constructed and then we present the results of our clustering and source separation algorithm when applied to this dataset.
I-A Synthetic multi-scale dataset
@@
  • Large scale ($w=2^{12}$). In this timescale, the signal alternates between two different noises to mimic the change in background noise due to day-night periodicity. The first noise is a white noise $\mathbf{x}_{1}(t)$ and the second one is a multifractal random noise $\mathbf{x}_{2}(t)$ which is a non-Gaussian noise with intermittency [75]. The large-scale signal is then $\mathbf{x}_{\text{large}}(t)=(1-\theta(t))\,\mathbf{x}_{1}(t)+\theta(t)\,\mathbf{x}_{2}(t)$ where $\theta(t)$ is equal to 0 on $[0,w]$ and $1$ on $w+\eta,2w-\eta]$ for small $\eta$, is $2w$-periodic and has a smooth junction at $t=w$ and at $t=2w$.
@@
The resulting dataset is a long time-series $\mathbf{x}=\mathbf{x}_{\text{large}}+\mathbf{x}_{\text{medium}}+\mathbf{x}_{\text{fine}}$. Fig. 24 shows one realization of this dataset along with the sources in three scales that compose it.
I-B Experimental details
While we are aware of the underlying sources that create this synthetic dataset, we choose to make an uninformed decision regarding the number of clusters per scale to mimic a more realistic setting. To this end, we use the same architecture and hyperparameters as the fVAE trained in the context of clustering and source separation for the Mars InSight seismic data. Specifically, the model architecture, training procedure, and optimization settings are directly borrowed from the Mars experiment to ensure comparability between real and synthetic datasets (see Sections V-B and V-C in the main text for details).
The only distinction is that, for the synthetic data, the timescales used for scattering covariance generation were selected to match the durations of the three source types in our multi-scale dataset. Specifically, we use the same window sizes described in the preceding Section I-A for the pyramidal scattering spectra computation. This ensures that the scattering features are well-adapted to the characteristic timescales of the synthetic sources.
I-C Synthetic data clustering